OpenStudy (anonymous):

please someone help me solve this equation step by step solve by elimination 5x+2y=-7 7x-3y=11

OpenStudy (anonymous):

1st multiply the top equation by 3 and the bottom by 2 to be able to eliminate the y components

OpenStudy (anonymous):

5x + 2y = -7 |x3| --> 15x+6y = -21 7x - 3y = 11 |x2| --> 14x-6y=22 you got x, then search for y

OpenStudy (anonymous):

what she said ha

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

after multiplying by 3 and 2 i add the 2 equations together to get x right

OpenStudy (anonymous):

1/29?

OpenStudy (anonymous):

yes thats correct for x now plug that back into one of the equations to solve for y

OpenStudy (anonymous):

ok this is where i keep getting stuck i can not for the life of me figure out how to do that

OpenStudy (anonymous):

what numbers do i multiply or whatever im suppose to do

OpenStudy (anonymous):

5x+2y=-7 change x with 1/29

OpenStudy (anonymous):

5(1/29) + 2y = -7 2y = -7 - (5/29) So. y = [-7-(5/9)] / 2

OpenStudy (anonymous):

so then is the answer 70/29

OpenStudy (anonymous):

cant be im so confused

OpenStudy (anonymous):

no, the answer is x = 1/29 and y = -3.59

OpenStudy (anonymous):

thats where i keep getting stuck, how u got that for y, will u put it in laimens terms

OpenStudy (anonymous):

5(1/29) + 2y = -7 2y = -7 - (5/29) So. y = [-7-(5/9)] / 2 --> = -3.59

OpenStudy (anonymous):

just follow what I did there and you should get it

OpenStudy (anonymous):

5(1/29) + 2y = -7 --> 5/29 +2y=-7 2y = -7 - (5/29) 2y = (-203-5)/29 2y = -208/29 y = -208/29 * 1/2 y = -104/29

OpenStudy (anonymous):

ok, but im still confused i am really really bad when it comes to this type of math it helps me to make notes how did you get the -203? i need explanations for every answer except the obvious ones

OpenStudy (anonymous):

make -7 in a fraction with denominator is 29, 29*-7=-203

OpenStudy (anonymous):

ok and then u subtract -5 from that

OpenStudy (anonymous):

okay I will walk you through it: So we have x= 1/29 now you know how we got that so plug 1/29 into 5x + 2y = -7 This becomes 5(1/29) + 2y = -7 -----> 5(1/29) = 5/29 So, subtract 5/29 from both sides: 2y = -7 - (5/29) -------Evaluate that Then divide by 2: Then y = [-7-(5/29)] / 2 and y = -3.59 So x = 1/29 and y = -3.59

OpenStudy (anonymous):

the solution is (1/29,-104/29)

OpenStudy (anonymous):

-104/29 is a fraction form from -3.59

OpenStudy (anonymous):

yeah you can leave it in fraction form its more accurate, I just evaluated it

OpenStudy (anonymous):

ok ty so much for you guys help and the trouble it took, thank you for being patient with me

OpenStudy (anonymous):

hehe not a problem

OpenStudy (anonymous):

no problem glad we could help!

OpenStudy (anonymous):

tina, you understood the procedure, right?

OpenStudy (anonymous):

I am seeing her on this place through out the day. Is something wrong Ms Holman

OpenStudy (anonymous):

need help solving this equation step by step soybean meal is 18% protein, cornmeal is 9% protein how many lbs of each should be mixed together to get 360lb mixture that is 15%